OpenStudy (anonymous):

3x - 4y = 6 6x - 8y = 10 The system shown has _____ solution(s).

OpenStudy (anonymous):

set them bot up in y=mx+b format

OpenStudy (anonymous):

well, you could also use elimination if you would like to get each variable.

OpenStudy (anonymous):

so if you multiplied the first equation by three you would get 6x-8y=12 and 6x-8y=10

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

x + y = k x - y = k The solution to the system shown is _____. (0,k) (2, 2) (k, 0)

OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is _____. (13/34, -15/34) (3/22, -7/22) (13/22, 1/22)

OpenStudy (anonymous):

^those are the wrong equations @brittnicoleee_